You recently accepted an assignment with Oxford Limited as a financial consultant. One of your first
assignments is the analysis of two proposed capital investment projects. Details of initial investments,
after-tax cash flows and average annual profits are represented below:
Year After-tax Cash flows
Initial investments
Cash flows
Average annual profits
1
2
3
4
Project A Project B
(R50 000)
R26 200
R12 200
R12 200
R5 000
R1 400
(R50 000)
R14 200
R14 200
R14 200
R14 200
R1 700
2.1. Which project has a shorter payback period? Motivate your answer by doing the relevant
calculations (answers expressed in years, months, and days)
1
Expert's answer
2020-12-08T09:56:12-0500

"=\frac{11600}{12200} = 0.95 \\

12 \times 0.95 = 11.40 \; months"

11 months and "(30 \times 0.4)" days

11 months and 12 days

Thus payback period = 2 years 11 month and 12 days


"= \frac{7400}{14200} = 0.52 \; years \\

12 \times 0.52 = 6.24 \; months"

6 months and "(30 \times 0.24)" days

6 months and 7 days

Thus payback period = 3 years 6 months and 7 days


Answer: project A has shorter payback period.